Calculate the heat capacity of a calorimeter if th Calculate heat capacity of a calorimeter
Answer #1


mass of benzoicacid = 5000 grams

HEAT OF COMBUSTION = 26.38 KJ/G

q released = 5000*26.38 = 131900 kj

heat capacity of calorimeter = q/DT = 131900/12.644

   = 10431.82 Kj/C

   = 1.0432*10^4 kj/c